# FreeDiff: Progressive Frequency Truncation for Image Editing with Diffusion Models

### v1

@@ 1 Introduction
Inspired by these observations and by examining the Fourier transform of the denoising network’s intermediate features, we hypothesize that the network indeed prioritizes the learning of frequency components in a manner that correlates with the noise level across timesteps. Thus, to edit a specific image, proper guidance should mainly focus on specific frequency bands. Our analysis of the diffusion model justify the common empirical findings and practices for different timesteps in image editing [6, 1, 21].
Building on this analysis, we propose a novel fine-tuning free approach to image editing, which performs frequency truncation progressively to refine the guidance towards the target region. Initial hyperparameter settings are provided for different editing types, while better editing results can usually be obtained by fine-tuning the hyperparameters based on the initial settings. Empirical results from extensive image experiments demonstrate that our frequency space refinement of guidance facilitates versatile and universal editing capabilities.
@@ 2.2 Inversion of diffusion
Editing real images requires a tractable path through diffusion, making inversion techniques essential. DDIM inversion[18] with deterministic settings, employing a small guidance scale, is a simple and representative technique that leads to an acceptable image reconstruction with minor errors. However, the restriction imposed by the small guidance scale often conflicts with the requirements of many editing tasks. To address this issue, null-text inversion (NTI)[12] optimizes null embeddings for different diffusion timesteps to capture specific image information, leading to a better reconstruction result and overcoming the limitations of small guidance scales. Meanwhile, AIDI[15] introduces an accelerated fixed-point inversion method, enabling the application of larger guidance scales in subsequent editing tasks. EFI[9] further enhances editing capabilities by adding noise of different scales to the image to obtain noisy latents, which are then corrected with network inference, a process akin to virtual inversion. For simplicity, we use DDIM Inversion with fixed-point iteration to invert latents.
@@ 3 Preliminaries
Score-based diffusion models The diffusion process, characterized by multi-level noise perturbations, can be formulated as the discretization of Stochastic Differential Equation (SDEs)[20] and can be reversed if the scores of all noise levels are known. Denote the encoded image latent as $x_{0}\in\mathbb{R}^{CHW}$. The objective for the denoising network $\epsilon_{\theta}$ is to learn the score $\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x_{0})$ for the perturbed data $x_{t}$ across all noise levels $\sigma_{t}$ at timestep $t$ [19, 7]. Different discretization schedules lead to different diffusion models[19, 7, 20]. With the DDIM schedule [18], the objective is:
@@
where $\delta(x_{0})$ is a Dirac delta function centered at $x_{0}$, $w(t)$ is a positive weighting function, typically set to 1[18, 7], and $\alpha_{t}\in(0,1]$ is the noise schedule coefficient that controls the noise level and decreases to nearly 00 as $t$ approaches $T$. For brevity, we denote the score of the perturbed data as $\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t})$.
Guidance To influence the generation process via conditional distributions, we focus on $\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|c)$, where the condition $c$ is the encoded embedding of the class labels, text prompt, etc. The conditional score can be expressed as [3]:
Classifier free guidance (CFG)[8] is one of the prevalent approaches, which formulates the class guidance as:
@@
where $\epsilon_{\theta}(x_{t},c)$ is the conditional score w.r.t. the encoded text prompt $c$, $\phi$ is the encoded embedding from a null (empty) string and $\epsilon_{\theta}(x_{t},\phi)$ is its corresponding unconditional score. We denote the guidance as $g_{t}$. It is common practice to enlarge the guidance by a scale $\gamma>1$ since $p^{\gamma}(c|x_{t})\propto p(x_{t}|c)/p(x_{t})$[8]. The overall score with CFG at timestep $t$ is:
$\hat{\epsilon}_{\theta}(x_{t})=\epsilon_{\theta}(x_{t},\phi)+\gamma g_{t}.$ (6)
DDIM Inversion DDIM [18] applies a non-Markovian formulation to the diffusion process, allowing both the forward and backward processes to be deterministic with a specific noise schedule. Deterministic DDIM samples $x_{t}$ from $x_{t+1}$ by:
@@
With the assumption that the network’s output should be close between small enough steps, that is, by substituting $\epsilon_{\theta}(x_{t+1})\approx\epsilon_{\theta}(x_{t})$, we can rewrite (7) to:
@@
To avoid the introduction of random noise during the inversion and generation processes and to make the analysis brief, in our study we adopt the deterministic DDIM formulation as in [18].
@@ 4.1 Diffusion prior from a frequency perspective
Figure 2: Visualized decoded intermediate features and Fourier transformed features from a generation process with SD v1.5[17], with the prompt “a lovely corgi running on a city street”. The first, second, third, and fourth rows display the decoded noisy latents $x_{t}$s, the decoded $\tilde{x}_{t:0}$s, the guidance $g_{t}$, and the power spectrum of $x_{0}$ with the SNR (signal-to-noise ratio) indicator (red box) at the corresponding timestep. The timestep is shown at the bottom. The SNR box indicates where the signal (image) to latent noise ratio is greater than 1, which suggest the frequency bands that the network has higher probability to successfully recover $x_{0}$ from $x_{t}$. Note that to show lower frequency components, the same power spectrum is normalized with lower truncated upper bound as $t$ decreases.
@@
We then visualize in Fig. 2 the decoded noisy latent $x_{t}$, the corresponding decoded intermediate image $\tilde{x}_{0|t}$, and the guidance $g_{t}$ (Eq. 5) from the frequency space using the 2D discrete Fourier transform (DFT). From the example, we observe decreasing noise from the decoded $x_{0}$ (first row), as expected from the denoising process. However, when observing the intermediate images, a distinct pattern emerges from $\tilde{x}_{0|t}$ (second row), revealing a process where finer details are progressively added across timesteps, which means higher frequency components are gradually incorporated. The difference between $x_{t}$ and $\tilde{x}_{0|t}$ is consistent with the nature of both $x_{0}$ and $\tilde{x}_{0|t}$ being the weighted sum of the noisy latent $x_{t}$ and the semantically guided $\hat{\epsilon}_{\theta}(x_{t})$, with $g_{t}$ having a larger weight in $\tilde{x}_{0|t}$. The frequency distribution of $g_{t}$ aligns with the visual transformations observed in $\tilde{x}_{0|t}$. Similar observations are obtained when inspecting other examples (see the Appendix).
@@
In the training process of the diffusion model, the noise introduced at timestep $t$ by (2) is additive white Gaussian noise (AWGN), characterized by uniform power across the frequency spectrum. Consider the image ultimately generated and displayed in Fig. 2, serving as one of the training samples with an energy spectrum following the power law. AWGN, with a constant energy spectrum of $\sigma_{t}^{2}=1-\alpha$ across all frequencies, is added to the encoded latent $x_{0}$. This summation results in varying signal-to-noise ratios (SNRs) across different frequency bands, with higher frequencies possibly being obfuscated by noise. The denoising network is tasked with learning to predict the original $x_{0}$ from the perturbed $x_{t}$ as accurately as possible. Yet, confined by the SNRs, the denoising network must primarily recover the low-frequency components at earlier timesteps (when the noise power is large), and progressively higher frequency components as the power of the AWGN decreases. In the fourth row of Fig. 2, the region with SNR $\geq 1$ is roughly outlined with a red box, serving as an indicator of which frequency bands the network could recover from the image at timestep $t$.
@@
where $\mathcal{E}$ denotes the encoder that transforms an image into its latent representation, and the summation performs channel-wise sum, with $C$ as the number of channels in the latent. $\mathcal{F}\{x\}(\omega)$ is the 2D frequency transform (2D DFT) of image $x$, and $\omega$ is the spatial frequency variable. In Fig. 3 we visualize $\mathcal{F}_{diff}(I_{src},I_{dirc-ed})$ and $\mathcal{F}_{diff}(I_{src},I_{attn-ed})$ for the source image $I_{src}$ and edited images based on direct editing $I_{dirc-ed}$ and attention-based editing $I_{attn-ed}$, which supports our hypothesis that a successful editing introduces less power in low-frequency components. More examples of different editing types with different attention-based methods supporting the hypothesis are provided in the Appendix.
@@ 4.2 Progressive frequency truncation
Figure 4: The pipeline of our proposed method. The progressive frequency truncation is only applied in the response period according to Alg. 4.2, while guidance outside the response period is set to zero.
@@
$\mathcal{M}_{t}^{S}=\mathcal{I}\big{(}\frac{\mathrm{abs}(\hat{g}_{t}-g_{t})}{\mathrm{abs}(g_{t})}>\kappa\big{)},\quad\tilde{g}_{t}=\hat{g}_{t}\circ\mathcal{M}_{t}^{S},$ (18)
@@
where $\kappa=0.6$ is the threshold, and $\eta_{0.8}(\tilde{g}_{t})$ denotes the $80\%$ percentile of $\tilde{g}_{t}$. We term this as $\eta$ truncation as opposed to the main spatial frequency truncation technique. Given these assumptions, the implementation of the progressive frequency truncation algorithm is detailed in Alg. 4.2 and visualized in Fig. 4.
@@
high-pass filter pairs and their upperbound timestep $\{(r_{t}^{H},r_{t}^{L},\tau_{i})\}$ Output: Refined guidance sequence $\{g_{t}^{\ast}\}$
@@
7:$\mathcal{M}_{t}^{S}=\mathcal{I}(\frac{\mathrm{abs}(\hat{g}_{t}-g_{t})}{\mathrm{abs}(g_{t})}>0.6)$
@@ 4.3 Application to editing
The categorization of editing types, when viewed from the perspective of frequency, varies significantly. Identity replacement, such as transforming a dog into a lion, is akin to object removal, with both focusing on modifications image textures, i.e., high spatial frequency (SF) information. In contrast, alterations in shape and pose correspond to adjustments in lower SF information. Changes in color and environmental (color) adjustments are associated with the lowest SF components, requiring specialized handling.
Given that the guidance amplitude for color in diffusion models is relatively low compared to that for objects, and since color information typically aligns with the lowest frequency components, merely applying frequency truncation proves ineffective for color editing. Instead, we propose a two-step process for color changes: first, generating a coarse mask for the object whose color is to be altered through frequency truncation on the guidance describing it at specific timesteps, for instance, "a white hat". Then, utilizing this mask solely to perform guidance truncation for the edit. The approach for changing the environment surrounding an object follows a similar methodology.
Empirically, we define hyperpermeter sets $(T_{st},T_{ed},r_{t}^{H},\tau_{i})$ for each editing type. In practical applications, $r_{t}^{L}$ is rarely employed due to its minimal impact in most scenarios. Detailed values for these hyperparameters are provided in the Appendix.
@@ 5 Experiments
We evaluate our method on on a variety of editing tasks and on a diverse set of images, most of which are sourced from the PIE benchmark[10]. For a fair comparison, all methods utilize the publicly available SD v1.5 [17] checkpoint, following the official implementations of the compared methods. Fixed-point iteration with $N=5$ is used for PNP, MasaCtrl and Our method to factor out the effect of inversion, except in cases where reconstruction fails with $N=5$. Null-text inversion is employed for P2P as it requires word-aligned source and target prompts to refine guidance with prompt-based attention operations. For inversion and re-generation with the target prompt, we perform the DDIM deterministic forward and backward sampling for 50 steps for all methods. The guidance scale is set to $7.5$ and the editing prompt remains the same across all methods unless specified otherwise.
@@ 5.1 Qualitative results
We present typical qualitative results demonstrating the comparative performance of our method against attention-based methods, including P2P, PNP, and MasaCtrl, as illustrated in Fig. 5. All showcased images are sourced from the PIE dataset. As evinced in Fig. 5, our method succeeds in performing precise editing across both rigid and non-rigid tasks, whereas P2P and PNP fail to handle non-rigid tasks, such as pose changes in row 3 and row 5. Our method achieves a satisfying balance between fulfilling the editing purpose and preserving the structure of the image. Moreover, our method is typically good at exerting subtle modifications, such as closing eyes and opening mouths, by leveraging our capability to adjust SF truncation to suit the editing task.
@@ 5.2 Quantitative results
Since our categorization of editing types differs from the existing PIE dataset[10], we have recategorized the dataset by applying fixed hyperparameters for each editing type to ensure a fair comparison. For consistency, detailed results and analysis are included in the Appendix.
@@ 7 Conclusion
Appendix
Appendix 0.A More intermediate features from generation process

### v2

@@ 1 Introduction
Inspired by these observations and by examining the Fourier transform of the denoising network’s intermediate features, we hypothesize that the network indeed prioritizes the learning of frequency components in a manner that correlates with the noise level across timesteps. Thus, to edit a specific image, proper guidance should mainly focus on specific frequency bands. Our analysis of the diffusion model justifies the common empirical findings and practices for different timesteps in image editing [6, 1, 21].
Building on this analysis, we propose a novel fine-tuning free approach to image editing, which performs frequency truncation progressively to refine the guidance towards the target region. Initial hyperparameter settings are provided for different editing types, whereas better editing results can usually be obtained by fine-tuning the hyperparameters based on the initial settings. Empirical results from extensive image experiments demonstrate that our frequency space refinement of guidance facilitates versatile and universal editing capabilities.
@@ 2.2 Inversion of diffusion
Editing real images requires a tractable path through diffusion, making inversion techniques essential. DDIM inversion[18] with deterministic settings, employing a small guidance scale, is a simple and representative technique that leads to an acceptable image reconstruction with minor errors. However, the restriction imposed by the small guidance scale often conflicts with the requirements of many editing tasks. To address this issue, null-text inversion (NTI)[12] optimizes null embeddings for different diffusion timesteps to capture specific image information, leading to a better reconstruction result and overcoming the limitations of small guidance scales. Meanwhile, AIDI[15] introduces an accelerated fixed-point inversion method, enabling the application of larger guidance scales in subsequent editing tasks. EFI[9] further enhances the editing capabilities by adding noise of different scales to the image to obtain noisy latents, which are then corrected with network inference, a process akin to virtual inversion. For simplicity, we use DDIM Inversion with fixed-point iteration to invert latents.
@@ 3 Preliminaries
For simplicity, we provide a brief overview here while offering a detailed version in the Appendix Sec.A, which explains all symbols and provides additional details.
Score-based diffusion models The diffusion process can be implemented as different discretization formulations of Stochastic Differential Equation (SDEs)[19, 7, 20, 18]. To avoid the introduction of random noise during the inversion and generation processes and to make the analysis brief, in our study we adopt the deterministic DDIM formulation as in [18]. The marginal distribution of noise perturbed latent is:
@@
where $\delta(x_{0})$ is a Dirac delta function centered at $x_{0}$ and $\alpha_{t}$ is the noise schedule coefficient. For brevity, we denote the score of the perturbed data as $\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t})$.
Guidance To control the generation, the guidance $g_{t}$ is commonly introduced by Classifier free guidance (CFG)[8] as the difference between and conditional score[3] $\epsilon_{\theta}(x_{t},c)$ and unconditional score $\epsilon_{\theta}(x_{t},\phi)$ as:
@@
The guidance $g_{t}$ is often enlarged by a factor $\gamma>1$.
DDIM Inversion Deterministic DDIM [18] inversion sample $x_{t}$ from $x_{t+1}$ by:
@@
and $\hat{\epsilon}_{\theta}(x_{t+1})$ can be approximated by $\hat{\epsilon}_{\theta}(x_{t})$:
@@ 4.1 Diffusion prior from a frequency perspective
Figure 2: Visualized decoded intermediate features and Fourier transformed features from a generation process with SD v1.5[17], with the prompt “a lovely corgi running on a city street”. The first, second, third, and fourth rows display the decoded noisy latents $x_{t}$s, the decoded $\tilde{x}_{t:0}$s, the guidance $g_{t}$, and the power spectrum of $x_{0}$ with the SNR (signal-to-noise ratio) indicator (red box) at the corresponding timestep. The timestep is shown at the bottom. The SNR box indicates where the signal (image) to latent noise ratio is greater than 1, which suggests the frequency bands that the network has higher probability to successfully recover $x_{0}$ from $x_{t}$. Note that to show lower frequency components, the same power spectrum is normalized with lower truncated upper bound as $t$ decreases.
@@
We then visualize in Fig. 2 the decoded noisy latent $x_{t}$, the corresponding decoded intermediate image $\tilde{x}_{0|t}$, and the guidance $g_{t}$ (Eq. 3) from the frequency space using the 2D discrete Fourier transform (DFT). From the example, we observe decreasing noise from the decoded $x_{0}$ (first row), as expected from the denoising process. However, when observing the intermediate images, a distinct pattern emerges from $\tilde{x}_{0|t}$ (second row), revealing a process where finer details are progressively added across timesteps, which means higher frequency components are gradually incorporated. The difference between $x_{t}$ and $\tilde{x}_{0|t}$ is consistent with the nature of both $x_{0}$ and $\tilde{x}_{0|t}$ being the weighted sum of the noisy latent $x_{t}$ and the semantically guided $\hat{\epsilon}_{\theta}(x_{t})$, with $g_{t}$ having a larger weight in $\tilde{x}_{0|t}$. The frequency distribution of $g_{t}$ aligns with the visual transformations observed in $\tilde{x}_{0|t}$. Similar observations are obtained when inspecting other examples (see the Appendix Sec.B).
@@
In the diffusion model training process, the noise introduced at timestep $t$ by (1) is additive white Gaussian noise (AWGN), characterized by uniform power across the frequency spectrum. Consider the image ultimately generated and displayed in Fig. 2, serving as one of the training samples with an energy spectrum following the power law. AWGN, with a constant energy spectrum of $\sigma_{t}^{2}=1-\alpha$ across all frequencies, is added to the encoded latent $x_{0}$. This summation results in varying signal-to-noise ratios (SNRs) across different frequency bands, with higher frequencies possibly being obfuscated by noise. The denoising network is tasked with learning to predict the original $x_{0}$ from the perturbed $x_{t}$ as accurately as possible. Yet, confined by the SNRs, the denoising network must primarily recover the low-frequency components at earlier timesteps (when the noise power is large), and progressively higher frequency components as the power of the AWGN decreases. In the fourth row of Fig. 2, the region with SNR $\geq 1$ is roughly outlined with a red box, serving as an indicator of which frequency bands the network could recover from the image at timestep $t$.
@@
where $\mathcal{E}$ denotes the encoder that transforms an image into its latent representation, and the summation performs channel-wise sum, with $C$ as the number of channels in the latent. $\mathcal{F}\{x\}(\omega)$ is the 2D frequency transform (2D DFT) of image $x$, and $\omega$ is the spatial frequency variable. In Fig. 3 we visualize $\mathcal{F}_{diff}(I_{src},I_{dirc-ed})$ and $\mathcal{F}_{diff}(I_{src},I_{attn-ed})$ for the source image $I_{src}$ and edited images based on direct editing $I_{dirc-ed}$ and attention-based editing $I_{attn-ed}$, which supports our hypothesis that a successful editing introduces less power in low-frequency components. More examples of different editing types with different attention-based methods supporting the hypothesis are provided in the Appendix Sec.C.
@@ 4.2 Progressive frequency truncation
Figure 4: The pipeline of our proposed method. The progressive frequency truncation is only applied in the response period according to Alg. 1, while guidance outside the response period is set to zero.
@@
$\mathcal{M}_{t}^{S}=\mathcal{I}\big{(}\frac{\mathrm{abs}(\hat{g}_{t}-g_{t})}{\mathrm{abs}(g_{t})}<\kappa\big{)},\quad\tilde{g}_{t}=\hat{g}_{t}\circ\mathcal{M}_{t}^{S},$ (15)
@@
where $\kappa=0.6$ is the threshold, and $\eta_{0.8}(\tilde{g}_{t})$ denotes the $80\%$ percentile of $\tilde{g}_{t}$. We term this as $\eta$ truncation, as opposed to the main spatial frequency truncation technique. Given these assumptions, the implementation of the progressive frequency truncation algorithm is detailed in Alg. 1 and visualized in Fig. 4.
@@
          high-pass filter pairs and their upperbound timestep $\{(r_{t}^{H},r_{t}^{L},\tau_{i})\}$
Output: Refined guidance sequence $\{g_{t}^{\ast}\}$
@@
7:$\mathcal{M}_{t}^{S}=\mathcal{I}(\frac{\mathrm{abs}(\hat{g}_{t}-g_{t})}{\mathrm{abs}(g_{t})}<0.6)$
@@ 4.3 Application to editing
The categorization of editing types, when viewed from the perspective of frequency, varies significantly. Identity replacement, such as transforming a dog into a lion, is akin to object removal, with both focusing on modifications of image textures, i.e., high spatial frequency (SF) information. In contrast, alterations in shape and pose correspond to adjustments in lower SF information. Changes in color and environmental (color) adjustments are associated with the lowest SF components, requiring specialized handling.
Given that the guidance amplitude for color in diffusion models is relatively low compared to that for objects, and since color information typically aligns with the lowest frequency components, merely applying frequency truncation proves ineffective for color editing. Instead, we propose a two-step process for color changes: first, generating a coarse mask for the object whose color is to be altered through frequency truncation on the guidance describing it at specific timesteps, for instance, "a white hat". Then, utilizing this mask solely to perform a guidance truncation for the edit. The approach for changing the environment surrounding an object follows a similar methodology.
Empirically, we define hyperpermeter sets $(T_{st},T_{ed},r_{t}^{H},\tau_{i})$ for each editing type. In practical applications, $r_{t}^{L}$ is rarely employed due to its minimal impact in most scenarios. Detailed values for these hyperparameters are provided in the Appendix Sec.D.
@@ 5 Experiments
We evaluate our method on on a variety of editing tasks and on a diverse set of images, most of which are sourced from the PIE benchmark[10]. For a fair comparison, all methods use the same SD v1.4 or SD v1.5[17] checkpoint, following the corresponding official implementations. Fixed-point iteration with $N=5$ is used for PNP, MasaCtrl and Our method to factor out the effect of inversion, except in cases where reconstruction fails with $N=5$. Null-text inversion is employed for P2P as it requires word-aligned source and target prompts to refine guidance with prompt-based attention operations. For inversion and re-generation with the target prompt, we perform the DDIM deterministic forward and backward sampling for 50 steps for all methods. The guidance scale is set to $7.5$ and the editing prompt remains the same across all methods unless specified otherwise.
@@ 5.1 Qualitative results
We present typical qualitative results that demonstrate the comparative performance of our method against attention-based methods, including P2P, PNP, and MasaCtrl, as illustrated in Fig. 5. All showcased images are sourced from the PIE dataset. As evinced in Fig. 5, our method succeeds in performing precise editing across both rigid and non-rigid tasks, whereas P2P and PNP fail to handle non-rigid tasks, such as pose changes in row 3 and row 5. Our method achieves a satisfying balance between fulfilling the editing purpose and preserving the structure of the image. Moreover, our method is typically good at exerting subtle modifications, such as closing eyes and opening mouths, by leveraging our capability to adjust the SF truncation to suit the editing task.
@@ 5.2 Quantitative results
Table 1: Quantitative results from partial dataset of PIE[10]
Ours P2P PNP MasaCtrl
CLIP Score (whole image) $\uparrow$ 25.5140 24.7521 25.4717 24.6614
Background LPIPS $\downarrow$ 11.14 11.83 15.01 13.97
We evaluate the CLIP score between the entire image and corresponding caption, and the LPIPS of the background region on the PIE dataset to assess semantic consistency and background preservation, respectively. However, we have identified two main issues within the PIE dataset, including incorrect categorization of image-editing type pairs and an ill-defined category, which hinders evaluation accuracy. To mitigate these issues, we selected a subset of approximately 200 images from PIE for a comprehensive evaluation, with the result shown in Tab. 1. Additionally, to illustrate the editing effect across subcategories, quantitative results are provided in Tab. 2. It is important to note that while the CLIP score serves as a metric for semantic consistency, it is not without limitations. Specifically, while the PNP method yields a higher CLIP score, its editing results are not comparable to those of the P2P method. For a detailed discussion of these issues and the rationale behind subset selection, please refer to the Appendix Sec.D.
Table 2: Quantitative results for corrected semantic categories, ’Cat’ denotes category and the number is consistent with the original PIE dataset.
Methods
[…]
MasaCtrl 24.57/.1661 24.83/.1001 25.58/.1810 26.92/.1043 25.01/.1190
PNP 25.30/.1733 26.03/.1053 25.77/.1997 26.92/.1293 26.45/.1328
P2P 24.78/.1340 25.11/.0889 24.02/.1768 27.14/.0835 25.76/.0941
Ours 24.97/.1253 26.49/.0798 24.17/.1428 27.47/.1341 25.74/.0972
@@ 7 Conclusion
Acknowledgements
This work was supported by a Strategic Research Grant from City University of Hong Kong (Project No.7005840).
Appendix of "FreeDiff: Progressive Frequency Truncation for Image Editing with Diffusion Models"
Appendix 0.A Preliminaries
Score-based diffusion models The diffusion process, characterized by multi-level noise perturbations, can be formulated as the discretization of Stochastic Differential Equation (SDEs) [20] and can be reversed if the scores of all noise levels are known. Different discretization formulations lead to different diffusion models[19, 7, 20, 18]. Denote the encoded image latent as $x_{0}\in\mathbb{R}^{CHW}$. The objective of the denoising network $\epsilon_{\theta}$ is to learn the score $\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x_{0})$ for the perturbed data $x_{t}$ across all noise levels $\sigma_{t}$ in the time step $t$[19, 7]:
where $w(t)$ is a positive weighting function, $\alpha_{t}\in(0,1]$ is the noise schedule coefficient that controls the noise level and decreases to nearly 00 as $t$ approaches $T$.
Guidance To influence the generation process via conditional distributions, we focus on $\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|c)$, where the condition $c$ is the encoded embedding of the class labels, text prompt, etc. The conditional score [3] can be expressed as:
Classifier free guidance [8] is often used in T2I diffusion models as in Eq.2 and Eq.3, where $\epsilon_{\theta}(x_{t},c)$ is the conditional score w.r.t. the encoded text prompt $c$, $\phi$ is the encoded embedding from a null (empty) string and $\epsilon_{\theta}(x_{t},\phi)$ is its corresponding unconditional score. It is common practice to enlarge the guidance by a scaling factor $\gamma>1$ since $p^{\gamma}(c|x_{t})\propto p(x_{t}|c)/p(x_{t})$, which equals to enhancing the posterior probability.
Appendix 0.B More intermediate features from generation process